# Test-Time Adaptation of Vision-Language Models for Open-Vocabulary Semantic Segmentation

### v1

@@ Appendix H Visualization of Layer-Wise Confidence Weights
In the main paper (Fig. 3), we presented the layer-wise confidence weights for the V20, Cityscapes, and COCO-Stuff datasets. Here, we extend the analysis by visualizing the weights for four additional datasets: V21, P59, P60, and COCO-Object in Fig. 6. As with the earlier results, we plot the mean and standard deviation of the learned weights across layers under various corruption types.
@@
Appendix I Visualization of Segmentation Maps
Figure 7 presents qualitative segmentation results on the v20 dataset, comparing predictions from the non-adaptive baseline, TENT, and our MLMP method. MLMP demonstrates stronger spatial consistency and fewer semantic errors, particularly in challenging regions where both the baseline and TENT struggle. The integration of intermediate-layer supervision appears especially beneficial for refining small object boundaries and correcting fine-grained details.
In Figures 816, we provide additional qualitative examples across a range of corruption types, using NaCLIP as the our OVSS backbone. These include both successful and failure cases under Gaussian noise, defocus blur, snow, and JPEG compression. Overall, the results illustrate the robustness of MLMP in mitigating noise-induced artifacts and improving prediction confidence, while also highlighting failure modes that warrant further investigation.
@@ Appendix I Visualization of Segmentation Maps
Appendix J Detailed Results of the Main Paper
@@
J.2 Comparison with Alternative Adaptation Methods
This section presents comprehensive tables with the full experimental results corresponding to those summarized in the main paper. Table 15 reports detailed results for the V21 dataset, Table 16 for P59, Table 17 for P60, and Table 18 for the Cityscapes dataset.
@@
CityScapes-C Average 21.63 21.64 21.60 13.45 23.02

### v2

@@ Appendix I Effect of Adaptation Iterations
In the main paper (Figure 3), we presented the layer-wise confidence weights for the V20, Cityscapes, and COCO-Stuff datasets. Here, we extend the analysis by visualizing the weights for four additional datasets: V21, P59, P60, and COCO-Object in Figure 5.1. As with the earlier results, we plot the mean and standard deviation of the learned weights across layers under various corruption types.
@@
Appendix L Visualization of Segmentation Maps
Figure 5.1 presents qualitative segmentation results on the v20 dataset, comparing predictions from the non-adaptive baseline, TENT, and our MLMP method. MLMP demonstrates stronger spatial consistency and fewer semantic errors, particularly in challenging regions where both the baseline and TENT struggle. The integration of intermediate-layer supervision appears especially beneficial for refining small object boundaries and correcting fine-grained details.
In Figures $5.1$, we provide additional qualitative examples across a range of corruption types, using NaCLIP as the our OVSS backbone. These include both successful and failure cases under Gaussian noise, defocus blur, snow, and JPEG compression. Overall, the results illustrate the robustness of MLMP in mitigating noise-induced artifacts and improving prediction confidence, while also highlighting failure modes that warrant further investigation.
@@
Appendix M Detailed Results of the Main Paper
@@
M.2 Comparison with Alternative Adaptation Methods
This section presents comprehensive tables with the full experimental results corresponding to those summarized in the main paper. Table 20 reports detailed results for the V21 dataset, Table 21 for P59, Table 22 for P60, and Table 23 for the Cityscapes dataset. In addition to the datasets reported above, we also include detailed results on the ACDC [acdc] dataset, which is specifically designed to capture real-world adverse conditions such as fog, night, rain, and snow. It is worth mentioning that in ACDC, approximately half of the adverse-condition images have corresponding reference (clean) views captured at nearly the same locations. While these reference images are captured at approximately the same location as the shifted ones, the scene content may differ (e.g., different vehicles or objects present) due to real-world variability. Following our main protocol, we perform test-time adaptation independently on both reference and adverse views, and report the mIoU for each condition in Table 24. The results show that MLMP consistently improves over TENT and the non-adapted baseline across all conditions, confirming its robustness to real, non-synthetic distribution shifts.
@@
[…]
Table 24: Detailed ACDC mIoU (reference vs. adverse views). All experiments use NACLIP as the OVSS model. MLMP consistently improves over TENT and the non-adapted baseline across all weather conditions.
OVSS: NACLIP Adaptation Method
Condition No Adapt. TENT MLMP
Fog (ref) 24.80 26.52 $\pm$0.03 31.28 $\pm$0.03
Fog 23.88 26.89 $\pm$0.04 33.33 $\pm$0.04
Night (ref) 24.95 26.54 $\pm$0.04 28.81 $\pm$0.10
Night 22.12 24.17 $\pm$0.00 24.76 $\pm$0.03
Rain (ref) 24.79 26.62 $\pm$0.00 31.10 $\pm$0.03
Rain 23.86 26.84 $\pm$0.04 32.44 $\pm$0.04
Snow (ref) 22.10 24.27 $\pm$0.04 28.22 $\pm$0.02
Snow 23.54 27.25 $\pm$0.05 30.59 $\pm$0.03
[…]